OpenStudy (mandii):

Solving Quadratic Equations with Complex Numbers x2+2x+9=0

jhonyy9 (jhonyy9):

x_1,_2=(-2+/- sqrt(4-36))/2=(-2+/- sqrt(-32))/2 =(-2+/- sqrt(-16*2))/2 = =(-2+/- 4isqrt(2))/2 =-1+/- 2isqrt2

OpenStudy (anonymous):

\[x^2+2x=-9\] \[(x+1)^2=-9+1=-8\] \[x+1=\pm\sqrt{-8}=\pm2\sqrt{2}i\] \[x=-1\pm2\sqrt{2}i\]